OpenStudy (anonymous):

Find the indicated derivative. Assume that all vector functions are differentiable. (expression in next post)

OpenStudy (anonymous):

\[\frac{ d }{ dt }(r_{1}(2t) + r _{2}(\frac{ 1 }{ t }) )\] note that (2t) and (1/t) are r AT (...) and this is not multiplication.

OpenStudy (dumbcow):

unless we know what r(t) is, all we can do is use chain rule \[\rightarrow 2r_1'(2t) -\frac{1}{t^{2}}r_2'(\frac{1}{t})\]

OpenStudy (anonymous):

That is all I wanted actually! But I don't see how using the chain rule means multiplying by the derivative of the input.

OpenStudy (anonymous):

Nevermind, I understand XD Thanks!

OpenStudy (dumbcow):

yw :)